OpenStudy (anonymous):

m^2- 1 7 ----m = ---- 2 2 Solve by completing the square

jhonyy9 (jhonyy9):

you dont believe me ?

jhonyy9 (jhonyy9):

2m*2 -m -7 = 0

OpenStudy (anonymous):

Is that equation typed correctly? I'm only asking because it'll be a messy answer, and completing the square is not the most efficient way to solve it.

OpenStudy (anonymous):

I don't understand it either...

OpenStudy (radar):

Collegegirl, the first fraction as shown is confusing. it presently can be shown as: m^3-m ------- 2 Is that what you want?

OpenStudy (radar):

Or is it m^2 -m/2 = 7/2

OpenStudy (anonymous):

Good question. Either way, completing the square is an awful way to solve it.